# Characterizing and Understanding HGNN Training on GPUs

### v3

@@ 4.1.3. Memory Pattern Analysis
DRAM Access Breakdown. Memory access latency is a critical factor contributing to performance degradation during model execution, and it is associated with high energy consumption. A comprehensive understanding of memory access patterns across various stages of HGNN training facilitates both performance and energy efficiency improvement. Fig. 6 illustrates the DRAM access breakdown of various models during the training process of HGNN on different datasets. From a training perspective, 7 the backward propagation generally incurs more memory access compared to the forward propagation. This phenomenon arises due to the reason that during the backward propagation, the model must access gradient data that matches the size of the dataset accessed in forward propagation. Additionally, the computation of gradients requires the accesses of numerous intermediate results preserved from the forward propagation. As presented in Fig. 6 (a), the backward propagation accounts for 55.38% of the total memory access during the entire training process, while the forward propagation averages only 42.87%.
Figure 6. DRAM access breakdown of HGNN training by phase: (a) Training; (b) Forward; (c) Backward.
@@
Memory Footprint Analysis. As the scale of real-world graph data continues to grow, memory utilization emerges as a pivotal factor influencing the scalability of models applied to large-scale datasets. Conducting an analysis of the memory footprint facilitates the optimization of model scalability. Fig. 7(a) illustrates the memory allocation in each execution stage during the training process of HGNN along the timeline. For the sake of convenience, only the specific cases of the HAN model on three different datasets are shown here, while the situations for the other two models are similar. Here, we compare the situations between the pure inference stage and the complete training stage to highlight the more urgent memory demands during the training process.
@@
In comparison to pure inference, the training process necessitates greater memory footprint, with the majority memory allocation lies on the NA stage. This distinction arises from inference’s singular execution of a forward propagation, which does not involve gradient computation and thereby obviates the storage of intermediate computational results essential for training during the forward propagation. Furthermore, the inference process does not require backward propagation, thus eliminating the associated memory allocation needed during training. And given that the numerous intermediate computation results requiring storage are associated with edges, the NA stage necessitates the most substantial memory allocation. Considering the case of DBLP dataset in Fig. 7(a), in comparison to the inference process, the total memory allocation during training increases by 1.92$\times$, with the NA stage constituting 66.63% of the total memory consumption.
@@
Cache Hit Rate Analysis. The cache hit rate effectively reflects data locality, thereby serving as a starting point for optimizing models from this perspective. Fig. 7(b) illustrates the differences in cache hit rates for forward and backward propagation across various stages of execution on different models of the HAN model. As shown in the figure, in nearly all scenarios, 10 the data locality during backward propagation tends to be lower compared to that observed during forward propagation. This discrepancy primarily arises because backward propagation involves not only accessing graph data but also retrieving stored gradient information and various intermediate variables, necessitating a broader range of memory accesses and leading to diminished data locality. For the FP, NA, and SF stages, the L2 cache hit rate during backward propagation is on average 23.71%, 3.57%, and 4.34% lower than during forward propagation, respectively.
@@ 4.1.4. Issue Stall Analysis
This section provides a detailed analysis of GPU instruction issue stalls encountered during the training process. The focus is on elucidating the temporal distribution and underlying causes of these stalls at each stage. Analyzing instruction issue stalls can reveal operational bottlenecks from an alternative perspective, thereby offering pathways for targeted optimization strategies.
Overall Profiling Results. Fig. 8(a) presents the ratio of stall time to execution time for different execution stages of the training process. From a comprehensive viewpoint, the prevalence of instruction issue stall across different stages throughout the training process is notable and merits careful consideration, averaging 33.21% across different stages. Fig. 8(c) illustrates the breakdown of instruction issue stall reasons for the main CUDA kernels invoked at each execution stage. It can be observed that 11 the vast majority of kernel stalls originate from memory dependency except for sgemm kernel, and the proportion of stall time attributable to memory dependency during the NA stage is the highest among all execution stages. This is because, during the training process of HGNN, other than sgemm, the kernels involve a large number of reads and writes to irregular graph data. These access patterns are usually highly irregular, making it difficult for the accessed data to be fully cached. Therefore, stalls caused by memory dependencies dominate. Furthermore, the NA stage, which predominantly accesses irregular graph data during aggregation, thereby resulting in the most significant instruction issue stalls attributable to memory dependency. As shown in Fig. 8(c), for kernels other than sgemm, memory dependency accounts for an average of 74.11% of the instruction issue stalls. In contrast, for the sgemm kernel, memory dependency accounts for an average of only 15.51%, with the dominant stalls being caused by execution dependency, which average 18.86%. Fig. 8(b) illustrates that the NA stage exhibits the highest ratio of stalls caused by memory dependency relative to total elapsed time across all execution stages during both forward and backward propagation.
Figure 8. Issue stall of HAN model on DBLP dataset: (a) The ratio of stall time to elapsed time.; (b) Memory dependency stall ratio to elapsed time; (c) Breakdown of issue stall reasons of different stages.
@@ 4.1.5. Exploring Metapath Changes
Increase in Length of Metapaths. Exploring the performance changes at different stages due to variations in metapath length is crucial, as longer metapaths can assist models in capturing more complex relationship patterns and context information over greater distances. Fig. 9(a) shows the execution time of different stages across varying metapath lengths. Obviously, 13 increase in the length of the metapath only significantly increase the execution time of NA stage, while the FP and SF stages are almost unaffected. The rationale behind this lies in the fact that the length increase of the metapath does not alter the vertex types or the number of semantic graphs, thereby maintaining the workloads on the FP and SF stages. However, elongating the metapath results in denser semantic graphs (Yan et al., 2022), characterized by a higher volume of edges. This directly translates into increased execution times during the NA stage. As depicted in Fig. 9(a), an increase in metapath length from 3 to 9 correlates with a 4.12$\times$ rise in execution time for the NA stage, while the FP and SF stages exhibit negligible change in execution times.
@@
Increase in Number of Metapaths. Exploring the impact of varying the number of metapaths on the execution performance of different stages is essential, as a greater number of metapaths can provide the model with richer semantic graph information, thereby enhancing model accuracy. Fig. 9(b) illustrates the impact of increasing the number of metapaths while keeping the metapath length constant on the performance at various execution stages. According to the figure, 14 increase in the number of metapaths leads to increased execution times across all stages of HGNN, with the NA stage showing the most noticeable growth trend. As the number of metapaths increases, so does the number of semantic graphs, resulting in a higher volume of vertices and edges to process. Consequently, each execution stage experiences a proportional increase in workload. Notably, the NA stage shows the most pronounced growth, primarily because the rise in edges outpaces that of nodes and even semantic graphs. As illustrated in Fig. 9(b), expanding the metapath count from 1 to 8 leads to execution time increases of 4.71$\times$ for FP, 9.34$\times$ for NA, and 1.40$\times$ for SF.
@@ 4.2. Mini-batch Training
As the size of real-world graph datasets continues to expand, mini-batch training has increasingly become the predominant method for model training, which leads to faster convergence and improved generalization performance compared to full-batch training. Aside from the additional mini-batch sampling stage, the execution behavior and characteristics are identical to those in full-batch training. Consequently, this section primarily focuses on profiling the mini-batch sampling stage. Unless otherwise specified, the batch size used for sampling is 256 for ACM, IMDB and DBLP datasets. Due to the large scale of the MAG dataset, its batch size is set to 1024. The sampling method employed is Random Walk Neighbor Sampler (Hamilton et al., 2017). Unless specified otherwise, the sampling process is executed on the CPU.
@@ 4.2.1. Execution Analysis
Execution Time Breakdown. Breaking down the execution time facilitates the identification of the primary execution processes during mini-batch training. Fig. 10(a) shows the breakdown of execution time across various execution stages. 15 Mini-batch sampling occupies the majority of the execution time in each epoch of mini-batch training, even exceeding the combined time of all other execution stages. This is due to mini-batch sampling in HGNN models requiring sampling operations across multiple semantic graphs not just one graph as in conventional GNNs, entailing traversal of intricate and irregular graph structures. Furthermore, this procedure is carried out on CPU, which is markedly time-intensive in contrast to the highly parallelized computational workload executed on GPU. As presented in Fig. 10(a), mini-batch sampling accounts for an average of 51.01% of the total time across different models and datasets. Especially for the HAN model on the MAG dataset, the time elpsed on sampling in each epoch accounts for as much as 88.92%.
@@
Characteristics Analysis of the Sampling Process. We conduct an in-depth analysis of the mini-batch sampling stage utilizing GPU-based sampling techniques, and present the performance metrics of the primary kernels in Table 5. The performance metrics detailed here pertain exclusively to the CUDA kernels specific to the mini-batch sampling stage, distinguishing them from those involved in the main execution stages.
Table 5. Profiling results of mini-batch sampling of HAN model on MAG datasets.
@@ 4.2.2. Exploring Metapath Changes
Increase in Length of Metapath. Fig. 11(a) and Fig. 11(b) illustrate the impact on the execution time of mini-batch sampling as metapath length varies. 17 For smaller datasets, the time required for mini-batch sampling exhibits minimal sensitivity to variations in metapath length. Conversely, large-scale datasets experience prolonged execution times as metapath length increases. Increasing the length of metapath augments the number of edges within the semantic graph, consequently elevating the number of edges sampled per batch. However, thread-level parallelism is utilized during sampling process. The additional sampling workload, facilitated by adequate hardware resources, does not significantly affect overall performance. Notably, sampling time increases only when the workload surpasses the hardware’s maximum parallel processing capacity. As shown in Fig. 11(a), compared to a metapath length of 3, the sampling time ratio remains around 1 as the metapath length increases on DBLP dataset. However, as depicted in Fig. 11(b), for larger datasets MAG, as the metapath length increases, the number of sampled edges in each batch grows more significantly, and the mini-batch sampling time also increases accordingly.
@@ 5. Multi-GPU Training
Distributed training, leveraging multiple GPUs for model training, is driven by the increasing complexity of real-world datasets and models, which can be implemented using either full-batch or mini-batch methods. In comparison to full-batch distributed training, mini-batch distributed training can markedly reduce the convergence time of the training process while preserving model accuracy. Furthermore, it demands fewer hardware resources, rendering it superior in terms of both performance and energy efficiency. Consequently, we focuse exclusively on the analysis of the mini-batch distributed training method in this section.
@@ 5.1.1. Performance Comparison with Different Number of GPUs
Fig. 12(a) illustrates the normalized execution time of distributed training using multiple GPUs compared to training on a single GPU. As depicted in the figure, 19 the performance improvement ratio in multi-GPU distributed training scenarios diverges from the ideal ratio. For smaller-scale datasets such as DBLP, increasing the number of devices may even result in a significant decline in overall training performance. This is mainly because an increase in the number of devices leads to competition for shared resources such as cache and bandwidth, thereby reducing performance. Relevant analysis will be provided in Section 5.2. As shown in Fig. 12(a), the overall execution time of two-GPUs and four-GPUs training are an average 1.37$\times$ and 8.69$\times$ compared to training on a single GPU for DBLP dataset. While linear improvements are evident in larger-scale datasets like MAG, scenarios involving two GPUs and four GPUs yield average performance enhancements of merely 1.66$\times$ and 2.13$\times$ respectively, falling short of the anticipated 2$\times$ and 4$\times$ improvements.
Figure 12. Overall resutls: (a) Normalized performance compared to trin on one GPU; (b) Execution time breakdown on DBLP dataset.
@@ 5.2. In-depth Analysis
5.2.1. CPU Resources Competition During Sampling
@@
To better highlight the decrease in sampling efficiency caused by CPU resource contention, we adopt an adaptive batch size approach here. Specifically, for each epoch, we shuffle the order of target vertices and evenly divide them into $num\_dst\_vertices/num\_gpus$ groups according to the number of GPUs. Each group is then sampled in parallel, and the complete batch after sampling is sent to a separate GPU for execution. We only present the relevant profiling results on the HAN model, with the results on other two models exhibit similar characteristics.
@@
As shown in Fig. 13(a), as the number of devices increases, the average number of edges sampled per batch for each GPU decreases. In theory, the average sampling time should also decrease, but in reality, the sampling time shows an increasing trend. Compared to a single GPU, in scenarios with two GPUs and four GPUs, the average time for mini-batch sampling is 1.29 $\times$ and 6.49 $\times$ that of a single GPU, respectively. 21 The primary factor limiting multi-process parallel sampling is the competition for shared CPU resources. As in Fig. 13(b), an increase in the number of context switches as the number of GPU increase indicates that due to the concurrent execution of multiple sampling processes, CPU cores become more finely divided in terms of time slice allocation, leading to increased overhead of context switching. The decrease in LLC hit rate indicates that contention for the shared cache by multiple sampling processes reduces the locality of cache data in each sampling process.
@@
Figure 13. CPU resources competition during mini-batch sampling: (a) Average nuumber of edges in one batch and related sampling time; (b) Number of CPU context switches and LLC hit rate.
5.2.2. Bandwidth Competition During Data Loading and Synchronization
Data loading and gradient synchronization are two critical steps in distributed training. 22 The factors limiting the performance of data loading and synchronization in multi-device scenarios are primarily the contention for shared bandwidth.

### v4

@@ 4.1.3. Memory Pattern Analysis
DRAM Access Breakdown. Memory access latency is a critical factor contributing to performance degradation during model execution, and it is associated with high energy consumption. A comprehensive understanding of memory access patterns across various stages of HGNN training facilitates both performance and energy efficiency improvement. Fig. 6 illustrates the DRAM access volume breakdown of various models during the training process of HGNN on different datasets. From a training perspective, 7 the backward propagation generally incurs more memory access compared to the forward propagation. This phenomenon arises due to the reason that during the backward propagation, the model must access gradient data that matches the size of the dataset accessed in forward propagation. Additionally, the computation of gradients requires the accesses of numerous intermediate results preserved from the forward propagation. As presented in Fig. 6 (a), the backward propagation accounts for 55.38% of the total memory access during the entire training process, while the forward propagation averages only 42.87%. The discrepancy in DRAM access volume between backward and forward propagation varies across different models and datasets, primarily due to the differing quantities of intermediate results stored during forward propagation, which are influenced by the complexity of model structure. This discrepancy is most pronounced in the HAN model on the ACM dataset, where the proportion of DRAM access for backward propagation is 2.32$\times$ greater than that for forward propagation.
Figure 6. Breakdown of DRAM access volume of HGNN training by phase: (a) Training; (b) Forward; (c) Backward.
@@
Memory Footprint Analysis. As the scale of real-world graph data continues to grow, memory utilization emerges as a pivotal factor influencing the scalability of models applied to large-scale datasets. Conducting an analysis of the memory footprint facilitates the optimization of model scalability. Fig. 7(a) illustrates the memory allocation in each execution stage during the training process of HGNN along the timeline. For the sake of convenience, only the specific cases of the HAN model on three different datasets are shown here, while the situations for the other two models are similar. Here, we compare the situations between the pure inference process and the complete training process to highlight the more urgent memory demands during the training process.
@@
In comparison to pure inference, the training process necessitates a greater memory footprint, with the majority of memory allocation lying on the NA stage. This distinction arises from inference’s singular execution of forward propagation, which does not involve gradient computation and thereby obviates the storage of intermediate computational results essential for training during the forward propagation. Furthermore, the inference process does not require backward propagation, thus eliminating the associated memory allocation needed during training. And given that the numerous intermediate computation results requiring storage are associated with edges, the NA stage necessitates the most substantial memory allocation. Considering the case of DBLP dataset in Fig. 7(a), in comparison to the inference process, the total memory allocation during training increases by 1.92$\times$, with the NA stage constituting 66.63% of the total memory consumption.
@@
Cache Hit Rate Analysis. The cache hit rate effectively reflects data locality, thereby serving as a starting point for optimizing models from this perspective. Fig. 7(b) illustrates the differences in cache hit rates for forward and backward propagation across various stages of execution on different models of the HAN model. As shown in the figure, in nearly all scenarios, 10 the data locality during backward propagation tends to be lower compared to that observed during forward propagation. This discrepancy primarily arises because backward propagation involves not only accessing graph data but also retrieving stored gradient information and various intermediate results, necessitating a broader range of memory accesses and longer reuse distance of data, leading to diminished data locality. For the FP, NA, and SF stages, the L2 cache hit rate during backward propagation is on average 23.71%, 3.57%, and 4.34% lower than during forward propagation, respectively.
@@ 4.1.4. Issue Stall Analysis
This section presents a comprehensive analysis of GPU instruction issue stalls observed throughout the training process. The emphasis is on elucidating both the temporal distribution and the underlying causes of these stalls at each stage of the process. Analyzing instruction issue stalls offers insights into operational bottlenecks from a distinct perspective, thereby facilitating the identification of targeted optimization strategies.
Overall Profiling Results. Fig. 8(a) presents the ratio of stall time to execution time for different execution stages of the training process. From a comprehensive viewpoint, the prevalence of instruction issue stall across different stages throughout the training process is notable and merits careful consideration, averaging 33.21% across different stages. Fig. 8(c) illustrates the breakdown of instruction issue stall reasons for the main CUDA kernels invoked at each execution stage. It can be observed that 11 the vast majority of kernel stalls originate from memory dependency except for sgemm kernel, and the proportion of stall time attributable to memory dependency during the NA stage is the highest among all execution stages. This is due to the fact that, during the training process of HGNNs, many kernels other than sgemm involve extensive read and write operations on irregular graph data, which is particularly evident in the graph-structure-related kernels such as SpMMCsr and SDDMMCoo employed in NA stage. These access patterns are usually highly irregular, making it difficult for the accessed data to be fully cached. Furthermore, elementwise operations in HGNN training may require accessing data distributed across non-contiguous memory locations, resulting in frequent and unpredictable memory access patterns. Therefore, stalls caused by memory dependencies dominate. As shown in Fig. 8(c), for kernels other than sgemm, memory dependency accounts for an average of 74.11% of the instruction issue stalls. In contrast, for the sgemm kernel, memory dependency accounts for an average of only 15.51%, with the dominant stalls being caused by execution dependency, which averages 18.86%. Fig. 8(b) illustrates that the NA stage exhibits the highest ratio of stalls caused by memory dependency relative to total elapsed time across all execution stages during both forward and backward propagation.
Figure 8. Issue stall of HAN model on DBLP dataset: (a) The ratio of stall time to elapsed time; (b) Memory dependency stall ratio to elapsed time; (c) Breakdown of issue stall reasons of different stages.
@@ 4.1.5. Exploring Metapath Changes
Increase in Length of Metapaths. Exploring the performance changes at different stages due to variations in metapath length is crucial, as longer metapaths can assist models in capturing more complex relationship patterns and context information over greater distances. Fig. 9(a) shows the execution time of different stages across varying metapath lengths. Obviously, 13 increase in the length of the metapath only significantly increases the execution time of NA stage, while the FP and SF stages are almost unaffected. The rationale behind this lies in the fact that the length increase of the metapath does not alter the vertex types or the number of semantic graphs, thereby maintaining the workloads on the FP and SF stages. However, elongating the metapath results in denser semantic graphs (Yan et al., 2022), characterized by a higher volume of edges. This directly translates into increased execution times during the NA stage. As depicted in Fig. 9(a), an increase in metapath length from 3 to 9 correlates with a 4.12$\times$ rise in execution time for the NA stage, while the FP and SF stages exhibit negligible change in execution times.
@@
Increase in Number of Metapaths. Exploring the impact of varying the number of metapaths on the execution performance of different stages is essential, as a greater number of metapaths can provide the model with richer semantic graph information, thereby enhancing model accuracy. Fig. 9(b) illustrates the impact of increasing the number of metapaths while keeping the metapath length constant on the performance at various execution stages. According to the figure, 14 increase in the number of metapaths leads to increased execution times across all stages of HGNN, with the NA stage showing the most noticeable growth trend. As the number of metapaths increases, so does the number of semantic graphs, resulting in a higher volume of vertices and edges to process. Consequently, each execution stage experiences a proportional increase in workload. Notably, the NA stage shows the most pronounced growth, primarily because the rise in edges outpaces that of vertices and even semantic graphs. As illustrated in Fig. 9(b), expanding the metapath count from 1 to 8 leads to execution time increases of 4.71$\times$ for FP, 9.34$\times$ for NA, and 1.40$\times$ for SF.
@@ 4.2. Mini-batch Training
As the size of real-world graph datasets continues to expand, mini-batch training has increasingly become the predominant method for model training, which leads to faster convergence and improved generalization performance compared to full-batch training. Aside from the additional mini-batch sampling phase, the execution behavior and characteristics are identical to those in full-batch training. Consequently, this section primarily focuses on profiling the mini-batch sampling phase. Unless otherwise specified, the batch size used for sampling is 256 for ACM, IMDB and DBLP datasets. Due to the larger scale of the MAG dataset, its batch size is set to 1024. The sampling method employed is Random Walk Neighbor Sampler (Hamilton et al., 2017). Unless specified otherwise, the sampling process is executed on the CPU.
@@ 4.2.1. Execution Analysis
Execution Time Breakdown. Breaking down the execution time facilitates the identification of the primary execution processes during mini-batch training. Fig. 10(a) shows the breakdown of execution time across various execution stages. 15 Mini-batch sampling occupies the majority of the execution time in each epoch of mini-batch training, even exceeding the combined time of all other execution stages. This is due to mini-batch sampling in HGNN models requiring sampling operations across multiple semantic graphs not just one graph as in conventional GNNs, entailing traversal of intricate and irregular graph structures. Furthermore, this procedure is carried out on CPU, which is markedly time-intensive in contrast to the highly parallelized computational workload executed on GPU. As presented in Fig. 10(a), mini-batch sampling accounts for an average of 51.01% of the total time across different models and datasets. Especially for the HAN model on the MAG dataset, the time elapsed on sampling in each epoch accounts for as much as 88.92%.
@@
Characteristics Analysis of the Sampling Process. We conduct an in-depth analysis of the mini-batch sampling phase utilizing GPU-based sampling techniques, and present the performance metrics of the primary kernels in Table 5. The performance metrics detailed here pertain exclusively to the CUDA kernels specific to the mini-batch sampling stage, distinguishing them from those involved in the main execution stages.
Table 5. Profiling results of mini-batch sampling of HAN model on MAG dataset.
@@ 4.2.2. Exploring Metapath Changes
Increase in Length of Metapath. Fig. 11(a) and Fig. 11(b) illustrate the impact on the execution time of mini-batch sampling as metapath length varies. 17 For smaller datasets, the time required for mini-batch sampling exhibits minimal sensitivity to variations in metapath length. Conversely, large-scale datasets experience prolonged execution times as metapath length increases. Increasing the length of metapath augments the number of edges within the semantic graph, consequently elevating the number of edges sampled per batch. However, thread-level parallelism is utilized during the sampling process. The additional sampling workload, facilitated by adequate hardware resources, does not significantly affect overall performance. Notably, sampling time increases only when the workload surpasses the hardware’s maximum parallel processing capacity. As shown in Fig. 11(a), compared to a metapath length of 3, the sampling time ratio remains around 1 as the metapath length increases on DBLP dataset. However, as depicted in Fig. 11(b), for larger datasets MAG, as the metapath length increases, the number of sampled edges in each batch grows more significantly, and the mini-batch sampling time also increases accordingly.
@@ 5. Multi-GPU Training
Distributed training, leveraging multiple GPUs for model training, is driven by the increasing complexity of real-world datasets and models, which can be implemented using either full-batch or mini-batch methods. In comparison to full-batch distributed training, mini-batch distributed training can markedly reduce the convergence time of the training process while preserving model accuracy. Furthermore, it demands fewer hardware resources, rendering it superior in terms of both performance and energy efficiency. Consequently, we focus exclusively on the analysis of the mini-batch distributed training method in this section.
@@ 5.1.1. Performance Comparison with Different Number of GPUs
Fig. 12(a) illustrates the normalized execution time of distributed training using multiple GPUs compared to training on a single GPU. As depicted in the figure, 19 the performance improvement ratio in multi-GPU distributed training scenarios diverges from the ideal ratio. For smaller-scale datasets such as DBLP, increasing the number of devices may even result in a significant decline in overall training performance. This is mainly because an increase in the number of devices leads to contention for shared resources such as cache and bandwidth, thereby reducing performance. Relevant analysis will be provided in Section 5.2. As shown in Fig. 12(a), the overall execution time of two-GPUs and four-GPUs training are an average of 1.37$\times$ and 8.69$\times$ compared to training on a single GPU for DBLP dataset. While linear improvements are evident in larger-scale datasets like MAG, scenarios involving two GPUs and four GPUs yield average performance enhancements of merely 1.66$\times$ and 2.13$\times$ respectively, falling short of the anticipated 2$\times$ and 4$\times$ improvements.
Figure 12. Overall results: (a) Normalized performance compared to training on one GPU; (b) Execution time breakdown on MAG dataset.
@@ 5.2. In-depth Analysis
5.2.1. CPU Resources Contention During Sampling
@@
To better highlight the decrease in sampling efficiency caused by CPU resource contention, we adopt an adaptive batch size approach here. Specifically, for each epoch, we shuffle the order of target vertices and evenly divide them into $num\_dst\_vertices/num\_gpus$ groups according to the number of GPUs. Each group is then sampled in parallel, and the complete batch after sampling is sent to a separate GPU for execution. We only present the relevant profiling results on the HAN model, with the results on the other two models exhibiting similar characteristics.
@@
As shown in Fig. 13(a), as the number of devices increases, the average number of edges sampled per batch for each GPU decreases. In theory, the average sampling time should also decrease, but in reality, the sampling time shows an increasing trend. Compared to a single GPU, in scenarios with two GPUs and four GPUs, the average time for mini-batch sampling is 1.29 $\times$ and 6.49 $\times$ that of a single GPU, respectively. 21 The primary factor limiting the performance of parallel sampling is the contention for shared CPU resources. As the number of GPUs increases, the concurrent execution of multiple sampling tasks in thread-level leads to finer partitioning of CPU cores in terms of time slice allocation, resulting in a higher frequency of context switching. As shown in Fig. 13(b), when the number of GPUs increases from 1 to 4, the number of CPU context switches increases by an average of 11.79$\times$. This process not only involves operations such as saving and restoring register states but also requires the operating system to schedule tasks based on their priorities. The cumulative delays introduced by these factors contribute to a significant decline in sampling performance. Moreover, the decrease in LLC hit rate suggests that contention among multiple sampling tasks for shared cache resources causes frequent data replacement within the cache, thereby increasing memory access latency and further exacerbating sampling delays.
@@
Figure 13. CPU resources contention during mini-batch sampling: (a) Average number of edges in one batch and related sampling time; (b) Number of CPU context switches and LLC hit rate.
5.2.2. Bandwidth Contention During Data Loading and Synchronization